### v2

@@ Abstract.
Dendric shifts are defined by combinatorial restrictions of the extensions of the words in their languages. This family generalizes well-known families of shifts such as Sturmian shifts, Arnoux-Rauzy shifts and codings of interval exchange transformations. It is known that any minimal dendric shift has a primitive $\mathcal{S}$-adic representation where the morphisms in $\mathcal{S}$ are positive tame automorphisms of the free group generated by the alphabet. In this paper we give an $\mathcal{S}$-adic characterization of this family by means of two finite graphs. As an application, we are able to decide whether a shift space generated by a uniformly recurrent morphic word is (eventually) dendric.
@@
${}^{\dagger}$upported by a FNRS fellowship.
@@ 1. Introduction
Dendric shifts are defined in terms of extension graphs that describe the left and right extensions of their factors. Extension graphs are bipartite graphs that can roughly be described as follows: if $u$ is a word in the language $\mathcal{L}(X)$ of the shift space $X$, one puts an edge between the left and right copies of letters $a$ and $b$ such that $aub$ is in $\mathcal{L}(X)$. A shift space is then said to be dendric if the extension graph of every word of its language is a tree. These shift spaces were initially defined through their languages under the name of tree sets [5] and were studied in a series of papers. They generalize classical families of shift spaces such as Sturmian shifts [30], Arnoux-Rauzy shifts [2], codings of regular interval exchange transformations [31, 1] (IET) or else shift spaces arising from the application of the Cassaigne multidimensional continued fraction algorithm [13] (MCF).
Minimal dendric shifts exhibit striking combinatorial [8, 10], algebraic [5, 7] and ergodic properties [4]. They for instance have factor complexity #$\#(\mathcal{L}(X)\cap\mathcal{A}^{n})=(\#\mathcal{A}-1)n+1$ [5] and topological rank #$\#\mathcal{A}$ [4], where $\mathcal{A}$ is the alphabet of the shift space. They also fall into the class of eventually dendric shift spaces, which are exactly those satisfying the regular bispecial condition [18]. This implies that the number of their ergodic measures is at most #$\#\mathcal{A}/2$ [16]. An important property for our work is that the derived shift of a minimal dendric shift is again a minimal dendric shift on the same alphabet, where derivation is here understood as derivation by return words. This allows to give $\mathcal{S}$-adic representations of such shift spaces [26], i.e., to define a set $\mathcal{S}$ of endomorphisms of the free monoid $\mathcal{A}^{*}$ and a sequence ${\boldsymbol{\sigma}}=(\sigma_{n})_{n\geq 0}\in\mathcal{S}^{\mathbb{N}}$, called an $\mathcal{S}$-adic representation, such that
@@
$\mathcal{S}$-adic representations are a classical tool that allows to study several properties of shift spaces such as the factor complexity [24, 19], the number of ergodic measures [9, 3, 12], the dimension group and topological rank [4] or yet the automorphism group [25]. In the case of minimal dendric shifts, the involved endomorphisms are particular tame automorphisms of the free group generated by the alphabet [8, 10]. This in particular allows to prove that minimal dendric shifts have topological rank equal to the cardinality of the alphabet and that ergodic measures are completely determined by the measures of the letter cylinders [4, 12].
In the case of the ternary alphabet $\{1,2,3\}$, we were able to give an $\mathcal{S}$-adic characterization of minimal dendric shifts [29]. We explicitly define an infinite set $\mathcal{S}_{3}$ of positive automorphisms of the free group $F_{3}$ and a labeled directed graph $\mathcal{G}$ with two vertices and labels in $\Sigma_{3}\mathcal{S}_{3}\Sigma_{3}$ (where $\Sigma_{3}$ is the symmetric group of $\{1,2,3\}$) such that a shift space over $\{1,2,3\}$ is a minimal dendric shift if and only if it has a primitive $\Sigma_{3}\mathcal{S}_{3}\Sigma_{3}$-adic representation labelling an infinite path in $\mathcal{G}$. We were then able to localize in $\mathcal{G}$ classical families of dendric shifts such as codings of regular interval exchanges or Arnoux-Rauzy shifts.
@@
Let $\mathcal{S}$ be a family of dendric return morphisms from $\mathcal{A}^{*}$ to $\mathcal{A}^{*}$ and let $X$ be a shift space having an $\mathcal{S}$-adic representation ${\boldsymbol{\sigma}}=(\sigma_{n})_{n\geq 0}$. Then $X$ is minimal dendric if and only if ${\boldsymbol{\sigma}}$ is primitive and labels infinite paths in the graphs $\mathcal{G}^{L}(\mathcal{S})$ and $\mathcal{G}^{R}(\mathcal{S})$.
Observe that, contrary to the ternary case, we are not able to explicitely define the set $\mathcal{S}$. We also consider two finite graphs instead of one. A key argument in the ternary case is that for all $N$, the $\mathcal{S}$-adic shift $X^{(N)}$ generated by $(\sigma_{n})_{n\geq N}$ is minimal and dendric, so that $\sigma_{N}$ preserves dendricity from $X^{(N+1)}$. The finiteness of $\mathcal{G}$ was obtained by defining an equivalence relation with finite index on the set of dendric shifts such that two shifts $X,Y$ are equivalent if and only if the same morphisms of $\Sigma_{3}\mathcal{S}_{3}\Sigma_{3}$ preserve dendricity from $X$ and from $Y$. The vertices of $\mathcal{G}$ are the two equivalent classes of this relation.
In the general case, while we are not able to explicitly define $\mathcal{S}$, we have enough information on the morphisms in $\mathcal{S}$ to similarly define an equivalence relation characterizing how dendricity is preserved. It is defined by means of two finite graphs describing the left and right extensions of infinite special factors of $X$ (see Section 4).
@@
Finally, in Section 8, we focus on the sub-family of interval exchanges and provide an $\mathcal{S}$-adic characterization (Theorem 72) using a subgraph of the graph obtained in the dendric case.
@@ 2.1. Words, languages and shift spaces
Let $\mathcal{A}$ be a finite alphabet of cardinality $d\geq 2$. Let us denote by $\varepsilon$ the empty word of the free monoid $\mathcal{A}^{*}$ (endowed with concatenation), by $\mathcal{A}^{+}$ the set of non-empty finite words on $\mathcal{A}$ and by $\mathcal{A}^{\mathbb{Z}}$ the set of bi-infinite words over $\mathcal{A}$. For a word $w=w_{1}\cdots w_{\ell}\in\mathcal{A}^{\ell}$, its length is denoted $|w|$ and equals $\ell$. We say that a word $u$ is a factor of a word $w$ if there exist words $p,s$ such that $w=pus$. If $p=\varepsilon$ (resp., $s=\varepsilon$) we say that $u$ is a prefix (resp., suffix) of $w$. For a word $u\in\mathcal{A}^{*}$ and a word $w\in\mathcal{A}^{*}\cup\mathcal{A}^{\mathbb{Z}}$, an index $j$ such that $w_{j}\cdots w_{j+|u|-1}=u$ is called an occurrence of $u$ in $w$. The number of occurrences of a word $u\in\mathcal{A}^{*}$ in a finite word $w$ is denoted by $|w|_{u}$.
The set $\mathcal{A}^{\mathbb{Z}}$ endowed with the product topology of the discrete topology on each copy of $\mathcal{A}$ is topologically a Cantor set. The shift map $S$ defined by $S\left((x_{n})_{n\in\mathbb{Z}}\right)=(x_{n+1})_{n\in\mathbb{Z}}$ is a homeomorphism of $\mathcal{A}^{\mathbb{Z}}$. A shift space is a pair $(X,S)$ where $X$ is a closed shift-invariant subset of some $\mathcal{A}^{\mathbb{Z}}$. It is thus a topological dynamical system. It is minimal if the only closed shift-invariant subset $Y\subset X$ are $\emptyset$ and $X$. Equivalently, $(X,S)$ is minimal if and only if the orbit of every $x\in X$ is dense in $X$. Usually we say that the set $X$ is itself a shift space.
@@ 2.2. Extension graphs and dendric shifts
The elements of $E_{F}^{L}(w)$, $E_{F}^{R}(w)$ and $E_{F}^{L}(w)$ are respectively called the left extensions, the right extensions and the bi-extensions of $w$ in $F$. If $X$ is a shift space over $\mathcal{A}$, we will use the terminology extensions in $X$ instead of extensions in $\mathcal{L}(X)$ and the index $\mathcal{L}(X)$ will be replaced by $X$ or even omitted if the context is clear. Observe that as $X\subset\mathcal{A}^{\mathbb{Z}}$, the set $E_{X}(w)$ completely determines $E_{X}^{L}(w)$ and $E_{X}^{R}(w)$. A word $w$ is said to be right special (resp., left special) if $\operatorname{Card}(E^{R}(w))\geq 2$ (resp., $\operatorname{Card}(E^{L}(w))\geq 2$). It is bispecial if it is both left and right special. The factor complexity of a shift space is completely governed by the extensions of its special factors. In particular, we have the following result.
@@
$=\sum_{w\in\mathcal{L}_{n}(X)}(\operatorname{Card}(E^{L}(w))-1)$
@@
A classical family of bispecial factors satisfying Equation (1) is made of the ordinary bispecial factors that are defined by $E(w)\subset(\{a\}\times\mathcal{A})\cup(\mathcal{A}\times\{b\})$ for some $(a,b)\in E(w)$. A larger family of bispecial factors also satisfying Equation (1) are the dendric bispecial factors defined below.
@@
Note that every non-bispecial word and every ordinary bispecial word is trivially dendric. In particular, the Arnoux-Rauzy shift spaces are dendric (recall that Arnoux-Rauzy shift spaces are the minimal shift spaces having exactly one left special factor $u_{n}$ and one right special factor $v_{n}$ of each length $n$ and such that $E^{L}(u_{n})=\mathcal{A}=E^{R}(v_{n})$; all bispecial factors of an Arnoux-Rauzy shift are ordinary). By Proposition 2, we deduce that any dendric shift has factor complexity $p_{X}(n)=(p_{X}(1)-1)n+1$ for every $n$.
@@
If $X$ is the Fibonacci shift generated by the Fibonacci word $\cdots01001010$. We have $\mathcal{L}_{3}(X)=\{001,010,100,101\}$. The extension graphs of the empty word and of the two letters 00 and $1$ are represented in Figure 1.
@@ 3. Extension graphs in morphic images
Let $\mathcal{A},\mathcal{B}$ be finite alphabets with cardinality at least 2. By a morphism $\sigma:\mathcal{A}^{*}\to\mathcal{B}^{*}$, we mean a monoid homomorphism (also called a substitution when $\mathcal{A}=\mathcal{B}$). We will always assume that $\mathcal{B}$ is minimal, i.e. each letter appear in the image of some word. A morphism is said to be non-erasing if the image of any letter is a non-empty word. We stress the fact that all morphisms are assumed to be non-erasing in the following. Using concatenation, we extend $\sigma$ to $\mathcal{A}^{\mathbb{Z}}$. In particular, if $X$ is a shift space over $\mathcal{A}$, the image of $X$ under $\sigma$ is the shift space
@@
In general, it can be difficult to deduce extension graphs of words in $\mathcal{L}(\sigma\cdot X)$ from those of $\mathcal{L}(X)$. In this paper, we will restrict ourselves to particular morphisms which offer some nice recognizability properties. They are defined using the notion of return words. Let $X$ be a shift space. If $w\in\mathcal{L}(X)$ is non-empty, a return word to $w$ in $X$ is a non-empty word $r$ such that $rw\in\mathcal{L}(X)$ and $rw$ contains exactly two occurrences of $w$, one as a prefix and one as a suffix.
@@
    Let $wu$ be a prefix of $w^{\prime}$. By definition, it is a prefix and a suffix of $\sigma(a)wu$ for all $a\in\mathcal{A}$. Moreover, since $w$ only has two occurrences in $\sigma(a)w$, $wu$ only has two occurrences in $\sigma(a)wu$ thus $\sigma$ is a return morphism for $wu$.
@@
    Since $\sigma$ is a return morphism for $w$, any occurrence of $w$ in $\sigma\cdot X$ is as a proper prefix of some $\sigma(a)w$ for $a\in\mathcal{A}$. However, for all $a\in\mathcal{A}$, $w^{\prime}$ is a prefix of $\sigma(a)w^{\prime}$ thus $ww^{\prime}_{|w|+1}$ is a prefix of $\sigma(a)w$. The only right extension of $w$ is then $w^{\prime}_{|w|+1}$.
@@
    Assume that $w^{\prime}$ only has one right extension $b$. Thus $w^{\prime}b$ is a prefix (and a suffix) of each $\sigma(a)w^{\prime}b$, $a\in\mathcal{A}$. As $w^{\prime}$ only has two occurrences in $\sigma(a)w^{\prime}$, the word $w^{\prime}b$ only has two occurrences in $\sigma(a)w^{\prime}b$, which proves that $\sigma$ is a return morphism for $w^{\prime}b$, a contradiction.
We will sometimes talk about a return morphism without specifying the word $w$ for which it is a return morphism. By default, this $w$ will be chosen of maximal length. Because of the previous proposition, most results are independent of this choice (or only depend in a minor way).
@@
  • If $u\in\mathcal{L}(Y)$ is such that $|u|_{w}=0$, then
@@
  • If $u\in\mathcal{L}(Y)$ is such that $|u|_{w}>0$, then there exists a unique triplet $(s,v,p)\in\mathcal{B}^{*}\times\mathcal{L}(X)\times w\mathcal{B}^{*}$ such that $u=s\sigma(v)p$ and for which the set $E_{X,s,p}(v)$ is not empty. The bi-extensions of $u$ are then governed by those of $v$ through the equation
@@
Let us prove the first case. For any $(b,c)\in E_{Y}(u)$, there exist a letter $a$ and a word $v\in\mathcal{L}(X)$ such that $buc$ is a factor of $\sigma(a)\sigma(v)$. We can moreover assume that the first occurrence of $buc$ begins in $\sigma(a)$. Since $w$ is not a factor of $u$ and $\sigma(v)$ is prefix comparable with $w$, i.e. $\sigma(v)$ is either a prefix of $w$ or has $w$ as a prefix, we deduce that $buc$ is a factor of $\sigma(a)w$. This proves that $(b,c)\in E_{\mathcal{L}(\sigma)}(u)$. By definition, $\mathcal{L}(\sigma)$ is included in $\mathcal{L}(Y)$ and we conclude that $\mathcal{E}_{Y}(u)=\mathcal{E}_{\mathcal{L}(\sigma)}(u)$.
@@
Under the assumptions of the previous result, the words $u$ that do not contain any occurrence of $w$ are the $\sigma$-initial factors of $Y$. As they and their extension graphs only depend on $\sigma$, we will sometimes talk about $\sigma$-initial factors without specifying $Y$. Note that a different convention for the choice of $w$ leads to a smaller set of $\sigma$-initial factors. However, by Proposition 6, the set of right special $\sigma$-initial factors is the same.
@@
Thus, any factor of $Y$ is either $\sigma$-initial or an extended image of some factor in $X$. Using Proposition 8, we directly know when the $\sigma$-initial factors of $Y$ are dendric. It only depends on $\sigma$ and not on $X$. We introduce the following definition.
@@
The morphism $\beta:0\mapsto 0,1\mapsto 01,2\mapsto 02,3\mapsto 032$ is a return morphism for 00. Its $\beta$-initial factors are $\varepsilon$, $1$, $2$, $3$ and $32$. Note that the only one which is right special is $\varepsilon$ and we can easily check that it is dendric thus $\beta$ is a dendric return morphism.
@@
Let $X$ be a shift space over $\mathcal{A}$, $\sigma:\mathcal{A}^{*}\to\mathcal{B}^{*}$ a return morphism for the word $w$, $Y=\sigma\cdot X$ and $u\in\mathcal{L}(Y)$ such that $|u|_{w}>0$. If $(s,v,p)$ is the triplet given by Proposition 8, then the extension graph of $u$ in $Y$ is the image of the graph $\mathcal{E}_{X,s,p}(v)$ by the morphisms $\varphi^{L}_{\sigma,s}$ acting on the left vertices and $\varphi^{R}_{\sigma,p}$ acting on the right vertices.
Remark that, as a consequence, $E^{L}_{Y}(u)\subset\varphi^{L}_{\sigma,s}(E^{L}_{X}(v))$ and $E^{R}_{Y}(u)\subset\varphi^{R}_{\sigma,p}(E^{R}_{X}(v))$. In particular, if $u$ is left (resp., right) special, $v$ is also left (resp., right) special and there exist two distinct letters $a,b\in\mathcal{A}$ and two distinct letters $a^{\prime},b^{\prime}\in\mathcal{B}$ such that

### v3

@@ Abstract.
Dendric shifts are defined by combinatorial restrictions of the extensions of the words in their languages. This family generalizes well-known families of shifts such as Sturmian shifts, Arnoux-Rauzy shifts and codings of interval exchange transformations. It is known that any minimal dendric shift has a primitive $\mathcal{S}$-adic representation where the morphisms in $\mathcal{S}$ are positive tame automorphisms of the free group generated by the alphabet. In this paper, we present an $\mathcal{S}$-adic characterization of this family using two finite graphs. As an application, we are able to decide whether a shift space generated by a uniformly recurrent morphic word is (eventually) dendric.
@@
The first author was supported by a FNRS fellowship.
@@ 1. Introduction
Dendric shifts are defined in terms of extension graphs that describe the left and right extensions of their factors. Extension graphs are bipartite graphs that can roughly be described as follows: if $u$ is a word in the language $\mathcal{L}(X)$ of the shift space $X$, one puts an edge between the left and right copies of letters $a$ and $b$ such that $aub$ is in $\mathcal{L}(X)$. A shift space is then said to be dendric if the extension graph of every word of its language is a tree. These shift spaces were initially defined through their languages under the name of tree sets [5] and were studied in a series of papers. They generalize classical families of shift spaces such as Sturmian shifts [30], Arnoux-Rauzy shifts [2], codings of regular interval exchange transformations (IET) [31, 1] , and shift spaces arising from the application of the Cassaigne multidimensional continued fraction algorithm (MCF) [13].
Minimal dendric shifts exhibit striking combinatorial [8, 10], algebraic [5, 7] , and ergodic properties [4]. For instance, they have factor complexity #$\#(\mathcal{L}(X)\cap\mathcal{A}^{n})=(\#\mathcal{A}-1)n+1$ [5] and topological rank #$\#\mathcal{A}$ [4], where $\mathcal{A}$ is the alphabet of the shift space. They also fall into the class of eventually dendric shift spaces, which are exactly those satisfying the regular bispecial condition [18]. This implies that the number of their ergodic measures is at most #$\#\mathcal{A}/2$ [16]. An important property for our work is that the derived shift of a minimal dendric shift is again a minimal dendric shift on the same alphabet, where derivation is here understood as derivation by return words. This allows to give $\mathfrak{S}$-adic representations of such shift spaces [26], i.e., to define a set $\mathfrak{S}$ of endomorphisms of the free monoid $\mathcal{A}^{*}$ and a sequence ${\boldsymbol{\sigma}}=(\sigma_{n})_{n\geq 0}\in\mathfrak{S}^{\mathbb{N}}$, called an $\mathfrak{S}$-adic representation, such that
@@
$\mathfrak{S}$-adic representations are a classical tool that allows to study several properties of shift spaces such as the factor complexity [24, 19], the number of ergodic measures [9, 3, 12], the dimension group and topological rank [4] or yet the automorphism group [25]. In the case of minimal dendric shifts, the involved endomorphisms are particular tame automorphisms of the free group generated by the alphabet [8, 10]. This in particular allows to prove that minimal dendric shifts have topological rank equal to the cardinality of the alphabet and that ergodic measures are completely determined by the measures of the letter cylinders [4, 12].
In the case of the ternary alphabet $\{1,2,3\}$, we were able to give an $\mathfrak{S}$-adic characterization of minimal dendric shifts [29]. We explicitly define an infinite set $\mathfrak{S}_{3}$ of positive automorphisms of the free group $F_{3}$ and a labeled directed graph $\mathcal{G}$ with two vertices and labels in $\Sigma_{3}\mathfrak{S}_{3}\Sigma_{3}$ (where $\Sigma_{3}$ is the symmetric group of $\{1,2,3\}$) such that a shift space over $\{1,2,3\}$ is a minimal dendric shift if and only if it has a primitive $\Sigma_{3}\mathfrak{S}_{3}\Sigma_{3}$-adic representation labelling an infinite path in $\mathcal{G}$. We were then able to localize in $\mathcal{G}$ classical families of dendric shifts such as codings of regular interval exchanges or Arnoux-Rauzy shifts.
@@
Let $\mathfrak{S}$ be a family of dendric return morphisms from $\mathcal{A}^{*}$ to $\mathcal{A}^{*}$ and let $X$ be a shift space having an $\mathfrak{S}$-adic representation ${\boldsymbol{\sigma}}=(\sigma_{n})_{n\geq 0}$. Then $X$ is minimal dendric if and only if ${\boldsymbol{\sigma}}$ is primitive and labels infinite paths in the graphs $\mathcal{G}^{L}(\mathfrak{S})$ and $\mathcal{G}^{R}(\mathfrak{S})$.
Observe that, contrary to the ternary case, we are not able to explicitely define the set $\mathfrak{S}$. We also consider two finite graphs instead of one. A key argument in the ternary case is that for all $N$, the $\mathfrak{S}$-adic shift $X^{(N)}$ generated by $(\sigma_{n})_{n\geq N}$ is minimal and dendric, so that $\sigma_{N}$ preserves dendricity from $X^{(N+1)}$. The finiteness of $\mathcal{G}$ was obtained by defining an equivalence relation with finite index on the set of dendric shifts such that two shifts $X,Y$ are equivalent if and only if the same morphisms of $\Sigma_{3}\mathfrak{S}_{3}\Sigma_{3}$ preserve dendricity from $X$ and from $Y$. The vertices of $\mathcal{G}$ are the two equivalent classes of this relation.
In the general case, while we are not able to explicitly define $\mathfrak{S}$, we have enough information on the morphisms in $\mathfrak{S}$ to similarly define an equivalence relation characterizing how dendricity is preserved. It is defined by means of two finite graphs describing the left and right extensions of infinite special factors of $X$ (see Section 4).
@@
Finally, in Section 8, we focus on the sub-family of interval exchanges and provide an $\mathfrak{S}$-adic characterization (Theorem 72) using a subgraph of the graph obtained in the dendric case.
@@ 2.1. Words, languages and shift spaces
Let $\mathcal{A}$ be a finite alphabet of cardinality $d\geq 2$ (note that some of the results do not work for unary alphabets). Let us denote by $\varepsilon$ the empty word of the free monoid $\mathcal{A}^{*}$ (endowed with concatenation), by $\mathcal{A}^{+}$ the set of non-empty finite words on $\mathcal{A}$ and by $\mathcal{A}^{\mathbb{Z}}$ the set of bi-infinite words over $\mathcal{A}$. For a word $w=w_{1}\cdots w_{\ell}\in\mathcal{A}^{\ell}$, its length is denoted $|w|$ and equals $\ell$. We say that a word $u$ is a factor of a word $w$ if there exist words $p,s$ such that $w=pus$. If $p=\varepsilon$ (resp., $s=\varepsilon$) we say that $u$ is a prefix (resp., suffix) of $w$. For a word $u\in\mathcal{A}^{*}$ and a word $w\in\mathcal{A}^{*}\cup\mathcal{A}^{\mathbb{Z}}$, an index $j$ such that $w_{j}\cdots w_{j+|u|-1}=u$ is called an occurrence of $u$ in $w$.
The set $\mathcal{A}^{\mathbb{Z}}$ endowed with the product topology of the discrete topology on each copy of $\mathcal{A}$ is topologically a Cantor set. The shift map $S$ defined by $S\left((x_{n})_{n\in\mathbb{Z}}\right)=(x_{n+1})_{n\in\mathbb{Z}}$ is a homeomorphism of $\mathcal{A}^{\mathbb{Z}}$. A shift space is a pair $(X,S)$ where $X$ is a closed shift-invariant subset of some $\mathcal{A}^{\mathbb{Z}}$. It is thus a topological dynamical system. It is minimal if the only closed shift-invariant subset $Y\subseteq X$ are $\emptyset$ and $X$. Equivalently, $(X,S)$ is minimal if and only if the orbit of every $x\in X$ is dense in $X$. Usually we say that the set $X$ is itself a shift space.
@@ 2.2. Extension graphs and dendric shifts
The elements of $E_{F}^{L}(w)$, $E_{F}^{R}(w)$ and $E_{F}^{L}(w)$ are respectively called the left extensions, the right extensions and the bi-extensions of $w$ in $F$. If $F$ is the language of a shift space $X$, we instead use the terminology extensions in $X$ and the index $F$ is replaced by $X$ or even omitted if the context is clear. Observe that as $X\subseteq\mathcal{A}^{\mathbb{Z}}$, the set $E_{X}(w)$ completely determines $E_{X}^{L}(w)$ and $E_{X}^{R}(w)$. A word $w$ is said to be right special (resp., left special) if $\operatorname{Card}(E^{R}(w))\geq 2$ (resp., $\operatorname{Card}(E^{L}(w))\geq 2$). It is bispecial if it is both left and right special. The factor complexity of a shift space is completely governed by the extensions of its special factors. In particular, we have the following result.
@@
$=\sum_{w\in\mathcal{L}_{n}(X)}(\operatorname{Card}(E^{L}(w))-1).$
@@
A classical family of factors satisfying Equation (1) is made of the ordinary factors that are defined by the existence of $(a,b)\in E(w)$ such that $E(w)\subseteq(\{a\}\times\mathcal{A})\cup(\mathcal{A}\times\{b\})$. A larger family of factors also satisfying Equation (1) are the dendric factors defined below.
@@
Note that every non-bispecial word and every ordinary bispecial word is trivially dendric. In particular, the Arnoux-Rauzy shift spaces are dendric (recall that Arnoux-Rauzy shift spaces are the minimal shift spaces having exactly one left special factor $u_{n}$ and one right special factor $v_{n}$ of each length $n$ and such that $E^{L}(u_{n})=\mathcal{A}=E^{R}(v_{n})$; all factors of an Arnoux-Rauzy shift are ordinary). By Proposition 2, we deduce that any dendric shift has factor complexity $p_{X}(n)=(p_{X}(1)-1)n+1$ for every $n$.
@@
Take a shift $X$ such that $\mathcal{L}_{3}(X)=\{001,010,100,101\}$ (for instance, the Fibonacci shift on $\{0,1\}$). The extension graphs of the empty word and of the two letters 00 and $1$ are represented in Figure 1.
@@ 3. Extension graphs in morphic images
Let $\mathcal{A},\mathcal{B}$ be finite alphabets with cardinality at least 2. By a morphism $\sigma:\mathcal{A}^{*}\to\mathcal{B}^{*}$, we mean a monoid homomorphism (also called a substitution when $\mathcal{A}=\mathcal{B}$). We will always assume that $\mathcal{B}$ is minimal, i.e. each letter appears in the image of some word. A morphism is said to be non-erasing if the image of any letter is a non-empty word. We stress the fact that all morphisms are assumed to be non-erasing in the following. Using concatenation, we extend $\sigma$ to $\mathcal{A}^{\mathbb{Z}}$. In particular, if $X$ is a shift space over $\mathcal{A}$, the image of $X$ under $\sigma$ is the shift space
@@
In general, it can be difficult to deduce extension graphs of words in $\mathcal{L}(\sigma\cdot X)$ from those of $\mathcal{L}(X)$. In this paper, we will restrict ourselves to particular morphisms which offer some nice recognizability properties in the sense that the pre-images of words are well understood. These morphisms are defined using the notion of return words. Let $X$ be a shift space. If $w\in\mathcal{L}(X)$ is non-empty, a return word to $w$ in $X$ is a non-empty word $r$ such that $rw\in\mathcal{L}(X)$ and $rw$ contains exactly two occurrences of $w$, one as a prefix and one as a suffix.
@@
    Let $wu$ be a prefix of $w^{\prime}$. By definition, it is a prefix and a suffix of $\sigma(a)wu$ for all $a\in\mathcal{A}$. Moreover, since $w$ has only two occurrences in $\sigma(a)w$, $wu$ has only two occurrences in $\sigma(a)wu$ thus $\sigma$ is a return morphism for $wu$.
@@
    Since $\sigma$ is a return morphism for $w$, $w$ only occurs in $\sigma\cdot X$ as a proper prefix of the words $\sigma(a)w$, $a\in\mathcal{A}$. However, for all $a\in\mathcal{A}$, $w^{\prime}$ is a prefix of $\sigma(a)w^{\prime}$ thus $ww^{\prime}_{|w|+1}$ is a prefix of $\sigma(a)w$. The only right extension of $w$ is then $w^{\prime}_{|w|+1}$.
@@
    Assume that $w^{\prime}$ has only one right extension $b$. Thus $w^{\prime}b$ is a prefix (and a suffix) of each $\sigma(a)w^{\prime}b$, $a\in\mathcal{A}$. As $w^{\prime}$ has only two occurrences in $\sigma(a)w^{\prime}$, the word $w^{\prime}b$ has only two occurrences in $\sigma(a)w^{\prime}b$, which proves that $\sigma$ is a return morphism for $w^{\prime}b$, a contradiction.
We will sometimes talk about a return morphism without specifying the word $w$ for which it is a return morphism. By default, this $w$ will be chosen of maximal length which is possible by the previous proposition. Because of the previous proposition, most results are independent of this choice (or only depend in a minor way).
@@
  • If $u\in\mathcal{L}(Y)$ does not contain any occurrence of $w$, then
@@
  • If $u\in\mathcal{L}(Y)$ contains an occurrence of $w$, then there exists a unique triplet $(s,v,p)\in\mathcal{B}^{*}\times\mathcal{L}(X)\times w\mathcal{B}^{*}$ such that $u=s\sigma(v)p$ and for which the set $E_{X,s,p}(v)$ is not empty. The bi-extensions of $u$ are then governed by those of $v$ through the equation
@@
Let us prove the first case. For any $(b,c)\in E_{Y}(u)$, there exist a letter $a$ and a word $v\in\mathcal{L}(X)$ such that $buc$ is a factor of $\sigma(a)\sigma(v)$. We can moreover assume that the first occurrence of $buc$ begins in $\sigma(a)$. Since $w$ is not a factor of $u$ and $\sigma(v)$ is prefix comparable with $w$, i.e. $\sigma(v)$ is either a prefix of $w$ or has $w$ as a prefix, we deduce that $buc$ is a factor of $\sigma(a)w$. This proves that $(b,c)\in E_{\mathcal{L}(\sigma)}(u)$. As every letter of $\mathcal{A}$ appears in $X$, $\mathcal{L}(\sigma)$ is included in $\mathcal{L}(Y)$ and we conclude that $\mathcal{E}_{Y}(u)=\mathcal{E}_{\mathcal{L}(\sigma)}(u)$.
@@
Under the assumptions of the previous result, the words $u$ that do not contain any occurrence of $w$ are the $\sigma$-initial factors of $Y$. As the set of these factors and their extension graphs only depend on $\sigma$ and not on $X$ and $Y$, we will sometimes talk about $\sigma$-initial factors without specifying $Y$. Note that a different convention for the choice of $w$ leads to a smaller set of $\sigma$-initial factors. However, by Proposition 6, the set of right special $\sigma$-initial factors is the same.
@@
Thus, any factor of $Y$ is either a $\sigma$-initial factor or an extended image of some factor in $X$. Using Proposition 8, we directly know when the $\sigma$-initial factors of $Y$ are dendric. It only depends on $\sigma$ and not on $X$. We introduce the following definition.
@@
The morphism $\beta:0\mapsto 0,1\mapsto 01,2\mapsto 02,3\mapsto 032$ is a return morphism for 00. Its $\beta$-initial factors are $\varepsilon$, $1$, $2$, $3$ and $32$. Note that the only one which is right special is $\varepsilon$ and we can easily check that it is dendric. Thus, $\beta$ is a dendric return morphism.
@@
Let $X$ be a shift space over $\mathcal{A}$, $\sigma:\mathcal{A}^{*}\to\mathcal{B}^{*}$ a return morphism for the word $w$, $Y=\sigma\cdot X$, and $u\in\mathcal{L}(Y)$ containing an occurrence of $w$. If $(s,v,p)$ is the triplet given by Proposition 8, then the extension graph of $u$ in $Y$ is the image of the graph $\mathcal{E}_{X,s,p}(v)$ by the morphisms $\varphi^{L}_{\sigma,s}$ acting on the left vertices and $\varphi^{R}_{\sigma,p}$ acting on the right vertices.
Remark that, as a consequence, $E^{L}_{Y}(u)\subseteq\varphi^{L}_{\sigma,s}(E^{L}_{X}(v))$ and $E^{R}_{Y}(u)\subseteq\varphi^{R}_{\sigma,p}(E^{R}_{X}(v))$. In particular, if $u$ is left (resp., right) special, $v$ is also left (resp., right) special and there exist two distinct letters $a,b\in\mathcal{A}$ and two distinct letters $a^{\prime},b^{\prime}\in\mathcal{B}$ such that